OpenStudy (anonymous):

Choose the equation of the line passing through the point (3, -2) and parallel to y = 3x - 6. y = 3x + 7 y = 3x - 11 y = -3x + 11 y = -3x - 12

OpenStudy (amistre64):

parallel lines differ at most by a constant; so the only that would change is that last term

OpenStudy (anonymous):

i think its the second one y = 3x - 11

OpenStudy (anonymous):

@amistre64 rkb is right

OpenStudy (amistre64):

give proof :)

OpenStudy (anonymous):

y - y1 = m(x-x1) y - (-2) = 3(x-3) perform that operation and you will get that

OpenStudy (anonymous):

okay -2 = 3(3) - 11 -2 = 9 -11 -2 = -2 i just plugged in the Points(:

OpenStudy (amistre64):

then we agree, 2nd option is good ;)

OpenStudy (amistre64):

the "how to" tends to be of more value than the final results ....